OpenStudy (anonymous):

How do i solve this to get the 2 answers of what x equals...x^2+x-12=0 I cant remember the tricks with the signs to figure out what makes 12 when you multiply them or am i completely wrong? The original equation is x(x+1)=12

OpenStudy (anonymous):

x^2 +4x -3x -12=0 x(x+4)-3(x+4)=0 (x+4)(x-3)=0 x=3,-4

OpenStudy (aroub):

u factor which becomes x^2 +x -12=0 , (x-3)(x+4)=0 so the answer may be 3 or -4 u have to check

OpenStudy (aroub):

or both

OpenStudy (anonymous):

Thank you! I get it now. I just couldn't remember.

OpenStudy (amistre64):

thats ok, happens to me alot these days; just cant remember things the way I used to :)